A cone made of paper is used to drink water from. It is made from circular paper with a radius of 12 cm, with a sector of 120° cut out. Calculate the area of paper needed to make the cone cup.


1
Expert's answer
2022-03-16T13:37:14-0400

"A=\pi r^2\frac{360-120}{360}=\pi 12^2*\frac{2}{3}=96\pi=301.59cm^2."